# Nathalie Gimenes PUC-Rio Tonghui Qi

### v1

@@ Abstract
How sellers choose reserve prices is central to auction theory, and the optimal reserve price depends on the seller’s risk attitude. Numerous studies have found that observed reserve prices lie below the optimal level implied by risk-neutral sellers, while the theoretical literature suggests that risk-averse sellers can rationalize these empirical findings. In this paper, we develop an econometric model of ascending auctions with a risk-averse seller under independent private values. We provide primitive conditions for the identification of the Arrow–Pratt measures of risk aversion and an estimator for these measures that is consistent and converges in distribution to a normal distribution at the parametric rate under standard regularity conditions. A Monte Carlo study demonstrates good finite-sample performance of the estimator, and we illustrate the approach using data from foreclosure real estate auctions in São Paulo.
@@ 1 Introduction
The reserve price is one of the seller’s primary tool for influencing expected revenue in an auction. Under the assumption of a risk-neutral seller, reserve prices play an essential role in the celebrated Revenue Equivalence Theorem of Myerson, (1981), while Riley and Samuelson, (1981) demonstrated that the optimal reserve is strictly positive, even with strong competition among bidders, across a range of auction formats. These insights form the benchmark that has guided much of the subsequent theoretical and empirical analysis of reserve-price behavior.
A number of empirical studies, however, have documented that observed reserve prices are well below the levels implied by the risk-neutral benchmark in a variety of settings. Examples include McAfee and Vincent, (1992), Paarsch, (1997), McAfee et al., (2002), Haile and Tamer, (2003), and Tang, (2011). One explanation is that sellers themselves are risk-averse, leading them to set lower reserves in order to increase the probability of a sale (Maskin and Riley, (1984), Matthews, (1987), Hu et al., (2010)).
Although the econometric literature on auctions has increasingly incorporated risk aversion into structural models, to date, the focus has been on the bidders in the context of first-price auctions (e.g., Lu and Perrigne, (2008), Guerre et al., (2009), Campo et al., (2011), Li et al., (2015), Zincenko, (2018), Grundl and Zhu, (2019)). We are not aware of any prior work that estimates the risk aversion of the seller despite its relevance in the theoretical literature.
In this paper, we propose an ascending auction model with risk-averse seller under the independent private value (IPV) framework. We provide primitive conditions, which parallel those assumed in the theoretical auction literature, under which Arrow-Pratt measures of risk aversion can be identified from the distribution of winning bid, reserve price, and seller’s valuation of the sale object. Subsequently, we show how the seller’s risk-aversion parameter in a parametric model, such as CARA or CRRA utility function, can be consistently estimated at the parametric rate with a limiting normal distribution. Our estimator is a two-step semiparametric estimator, as the risk parameter is identified by the first-order condition the seller uses for setting an optimal reserve, which involves the quantile of the bidder’s valuation that we flexibly estimate in the first step.
We perform a Monte Carlo study and find that our estimator behaves as expected from the theory. We then apply our methodology to real estate foreclosure auction data in São Paulo as an illustration. In this application, the sellers are lenders (banks and private companies) and the Court of Justice of the State of São Paulo. Our finding indicates that sellers are risk-averse, and the model with a risk-averse seller fits the data better than one with a risk-neutral seller.
Beyond explaining low reserve prices, knowing the seller’s risk attitude has policy implications. Auction design matters more when the seller is risk averse. For examples, the Revenue Equivalence Theorem does not hold with a risk-averse seller, and while some optimal auction results in Myerson, (1981) extend to risk-averse sellers, they may do so only in some cases (Sundararajan and Yan, (2020)). The reason for why sellers may be risk-averse depends on applications. In ours, it is straightforward to justify why sellers are averse to not being able to sell foreclosed properties. Firstly, there may be direct costs for holding on to such properties (e.g., property taxes, insurance, security). The value of foreclosure homes also tend to devalue over time. Moreover, from the accounting perspective, foreclosure assets are typically viewed as non-performing loans and sellers may have directives to sell them within a certain time frame [^1] This is indeed a guideline of some central banks, for example, see this report by the ECB https://www.bankingsupervision.europa.eu/ecb/pub/pdf/guidance_on_npl.en.pdf.
There are two steps to estimating our model. The first is the estimation of the bidders’ valuations distribution from bids data. The second is on the estimation of the risk parameter. The first stage of our analysis builds on the large literature on the identification and estimation of bidders’ valuations distribution from bids data, which starts with the parametric model by Paarsch, (1992) and grows with Donald and Paarsch (1993,1996), Laffont et al., (1995), Athey and Levin, (2001), Hirano and Porter, (2003), Li and Zheng, (2012) amongst others. The parametric models, however, can be computationally demanding to estimate and is subjected to model misspecification. The nonparametric approach, which begins with the seminal work of Guerre et al., (2000), offers an alternative that have been developed in a number of contexts, inter alia, see Athey and Haile, (2002), Lu and Perrigne, (2008), Krasnokutskaya, (2011), Marmer and Shneyerov, (2012), Hubbard et al., (2012), Campo et al., (2011), Marmer et al., (2013), Enache and Florens, (2017), Liu and Luo, (2017), Luo and Wan, (2018), Zincenko, (2018) and Ma et al., (2019). The purely nonparametric approach is also not without issues, as it suffers from the curse-of-dimensionality. Given the well-known trade-offs between parametric and nonparametric modelling, our work adopts the augmented quantile regression (AQR) approach of Gimenes and Guerre, (2022) (hereafter GG22) that uses a linear quantile specification to model the distribution of the bidder’s valuation. [^2] GG22 also proposed a more flexible version of the AQR where linear quantile is replaced by nonparametric additive functions.
AQR is a novel way to estimate quantiles by minimizing an integrated version of the classic Koenker and Bassett, (1978) objective function, combined with kernel smoothing and polynomial approximation of the quantile function (cf. Fan (1992,1993)). The local polynomial formulation is appealing to our application as it improves boundary properties and simultaneously delivers estimates of the quantile function and its derivatives; the latters enter the seller’s first-order condition that we use to estimate the seller’s risk aversion parameter. Our AQR estimator of bidder valuation quantiles in ascending auctions in fact targets the same object of interest as in Gimenes, (2017), who estimated it using the classic Koenker and Bassett objective function without local-polynomial features. It should be emphasized that our work does not contribute to the methodological development of the AQR estimator itself. However, our paper goes some way in expanding the range of its applications beyond the results stated in GG22, as the authors’ focus was on estimating bidder valuation quantiles from first-price bids.
In particular, we provide pointwise properties of the AQR quantile estimator and its inverse. The former was not given in GG22, although they provided the uniform convergence rate of the quantile estimator. They did not consider the latter, which serves as a flexible estimator of a conditional CDF and is of independent interest. We present these results in general form and apply them to an ascending auction model. Moreover, our analysis of the seller’s risk-aversion measure, which had not been estimated previously, makes use of Theorem 4 in GG22 that establishes asymptotic normality for integral functionals of AQR estimators. GG22 motivate the study of such functionals with examples that have closed-form expressions in the quantile functions (expected revenue (Li et al., (2003)) and bidder’s risk aversion parameter in a first-price auction (Guerre et al., (2009))). We show that such functionals can also be obtained when the AQR estimator is used to estimate nuisance functions in a general two-step estimation problem.
To estimate the seller’s risk-aversion measure in the second step, we assume that the seller’s utility function is known parametrically. Our risk aversion estimator satisfies a first-order condition that depends on the derivative and inverse of the bidder valuation quantile function, which we estimate in the first stage using the AQR method. To derive its limiting distribution, we take the pathwise differentiation approach to capture the contribution of the first-stage estimator by computing the functional derivative of the first-order condition with respect to the nuisance functions. This gives us a Riesz representer that, together with a stochastic equicontinuity property, enables us to express the risk-aversion parameter as an integral functional of the quantile of winning bids and its derivative. This particular representation allows us to invoke Theorem 4 of GG22 to establish a CLT. Our analysis otherwise follows the general procedure to obtain the limiting distribution of a semiparametric extremum estimator, for example, see Newey, (1994), Chen et al., (2003), and Ichimura and Lee, (2010).
@@
The rest of the paper is organized as follows. Section 2 introduces the model, establishes identification of bidder valuations, and characterizes optimal reserve-price behavior under risk aversion. Section 3 develops the estimation methodology and the asymptotic properties of the AQR estimators for observed bids and bidder valuations. Section 4 presents the estimation of the risk-aversion parameter and derives its asymptotic properties. Section 5 reports a Monte Carlo study of the estimators. Section 6 provides an empirical application to real estate auctions in São Paulo. Section 7 concludes. Proofs of the main results are collected in the Appendix.
@@ 2 Model and identification
Consider an ascending auction of an indivisible object with $I\geq 2$ bidders. Bidders can raise prices continuously and without cost until only one bidder remains. The object is thus sold to the highest bidder for the price of his last bid, provided that it is at least as high as the reservation price. We assume an independent private values (IPV) environment, where each bidder knows only their own valuation and values are independently drawn across bidders.
Let the auctioned object have observable characteristics $X\in\mathbb{R}^{D}$. Bidder $i$’s valuation of the object is $V_{i}$, taking value in $\mathcal{V}=\left[\underline{v},\overline{v}\right]$. For notational simplicity, we assume $\mathcal{V}$ to be independent of $X$. We denote the conditional CDF of $V_{i}$ by $F\left(\cdot|X\right)$. To facilitate readers, as it may be instructive to compare our assumptions and results with GG22, we try to adopt their notations and termiologies when possible in this and the next section.
@@ 2.1 Bidders’ behavior
The winning bid, denoted by $B=V_{i}^{I-1:I}\in\mathcal{V}$, is the $\left(I-1\right)$-th order statistic among the $I$ i.i.d. private values $\left\{V_{i}\right\}_{i=1}^{I}$. This is the same equilibrium play for ascending auctions as used in Aradillas-López et al., (2013) and Gimenes, (2017). We denote the conditional CDF of $B$ by $G\left(\cdot|X\right)$.
@@
Assumption M1.
(i) $F\left(\cdot|X\right)$ is differentiable and strictly increasing almost surely on $\mathcal{V}$;
Our paper takes a quantile approach to model the bidder’s private value distribution. Let the $\alpha$-quantile of the valuation distribution be denoted by $V\left(\alpha|X\right)=F^{-1}\left(\alpha|X\right)$ for $\alpha\in\left[0,1\right]$. As done in Gimenes, (2017) and GG22, we take $V_{i}=V\left(A_{i}|X\right)$ where $A_{i}\sim Uni\left[0,1\right]$ can be viewed as the private rank of the bidder, which is independent of $X$ and other bidders’ ranks.
@@
The quantile functions of the winning bid and the private valuation are linked through a one-to-one relationship. Proposition 1, which maps the quantile function of the private valuation from the quantile function of the winning bid, is the central identification argument on the bidder’s side (Gimenes, (2017)).
@@ 2.2 Seller’s behavior
The seller, whose value of the sale object is $W$, sets a reserve price, $R$, optimally to maximize her expected utility. Specifically, if the seller has utility function $U\left(\cdot\right)$, the expected utility from setting reserve price to be $r\in\mathcal{V}$ is:
@@
We assume $R=\arg\max_{{}_{r\in\mathcal{V}}}\widetilde{\Pi}\left(r,X,W\right)$, whose existence and uniqueness are guaranteed under the conditions of Assumption M2 given below.
Complementary to writing the bidders’ bids in terms of private ranks, the expected revenue can be equivalently expressed as a function of the screening level instead of the reserve price. For a screening level $\alpha$ that takes value in $\left[0,1\right]$, let us define
@@ 3 Augmented quantile regression
Given data on winning bids and auction characteristics, this section proposes estimators for the quantile function of private values and related functions. We give the pointwise asymptotic properties of the quantile function estimator and the convergence rates for the estimators of the derivatives and inverse of the quantile function. The convergence rates are needed for deriving the large sample properties of the risk-aversion estimator in Section 4.
Our estimation strategy proceeds in two steps. We first estimate the quantile function of the winning bids. Then, we apply the identity linking bid and value quantiles given in Proposition 1. The first step follows what GG22 calls the augmented quantile regression (AQR) approach. Estimators for other functions of interest can be obtained subsequently. Note that seller’s specific variables, namely reserve price and outside value for the sale object, are not used for quantile estimation.

### v2

@@ Abstract
This paper shows how to identify and estimate the seller’s risk parameter in an ascending auction. We consider a semiparametric model where the seller has a parametric utility function (such as CARA or CRRA) and the distribution of bidder valuations is modeled flexibly. We provide primitive conditions under which the risk parameter is identified and show that it can be consistently estimated with an asymptotically normal limiting distribution under standard regularity conditions. A Monte Carlo study demonstrates good finite-sample performance of the proposed estimator. We apply our approach to foreclosure real estate auction data from São Paulo. We find evidence that sellers are risk-averse, which leads to a much better fit to the data than a model with risk-neutral sellers, which would substantially underpredict the reserve price relative to what is observed.
@@ 1 Introduction
Knowing the seller’s risk attitude is essential for auction design and policy analysis, as seller risk preferences affect supply-side welfare and equilibrium outcomes. The theoretical literature has shown that many canonical results derived under risk neutrality no longer hold when sellers are risk averse. For example, the Revenue Equivalence Theorem does not hold with a risk-averse seller, and while some optimal auction results in Myerson, (1981) extend to risk-averse sellers, they do so only in some cases (Sundararajan and Yan, (2020)). More generally, seller risk aversion affects key design choices, including reserve prices and auction formats (Waehrer et al., (1998), Moreno and Wooders, (2017)). These insights highlight why policymakers seeking to predict auction outcomes and design effective auction mechanisms would want to learn about seller risk attitudes.
Prior to our study, to the best of our knowledge, there are no results on the identification of seller risk aversion or on how to estimate it in any auction format. Our paper addresses this gap in the literature. We study a semiparametric model of an ascending auction in which the seller has a parametric utility function (such as CARA or CRRA), while bidder valuations and independent and satisfy a linear quantile specification. We show that the risk-aversion parameter is identified under primitive conditions. We then propose an estimator for the risk-aversion parameter and show that our estimator is consistent and asymptotically normal under standard regularity conditions.
Using data on winning bids, reserve price, and value of the auctioned object, we can identify components of the first-order condition (FOC) that characterizes the seller’s optimal reserve price chosen to maximize expected revenue. Under interpretable conditions that parallel those assumed in the theoretical auction literature, part (i) of our Proposition 2 establishes existence and uniqueness of the solution to the FOC, while part (ii) shows that the optimal reserve price is decreasing in the seller’s Arrow–Pratt measure of risk aversion. We use the latter monotonic relationship between the optimal reserve price and risk aversion to establish identification of parametric risk preferences that include CARA and CRRA utility functions as special cases.
Our estimator of the risk-aversion parameter is a two-step parameter. In the first stage, nuisance functions, which are components of the expected revenue related to the distribution of bidder valuations, are estimated. The risk-aversion estimator is then defined by forcing the empirical FOC in the second stage. The nuisance functions consist of the derivative of the quantile function and the inverse of the quantile function. While only nonparametric restrictions on the bidder valuation distribution are required for identification, we further assume that the valuation quantile is linear in covariates for the purpose of practical estimation. This modeling choice represents a compromise between fully nonparametric and parametric alternatives, since the linear quantile specification allows for a flexible representation of conditional valuation distributions without suffering from the curse of dimensionality. In principle, other specifications for the distribution of bidder valuations can be adopted.
The use of quantiles in the econometrics of auctions is not new [^1] An early insight can be found in Haile et al., (2003). Nonparametric estimation of quantiles is used by Marmer and Shneyerov, (2012) and Guerre and Sabbah, (2012) in a first-price auction model, for example., but the development of a quantile regression (QR) estimator in this context is relatively recent. Methodologically, our approach builds on the estimators used in Gimenes, (2017) and Gimenes and Guerre (2022, GG22 hereafter). Their strategy exploits auction-specific relations between quantiles of the bidder’s valuation and their bids, in which they use bid data to estimate the QR for bids and transform it into a QR estimator for valuations. Specifically, Gimenes, (2017) applies this strategy to estimate ascending auctions with the classic Koenker-Bassett QR estimator, while GG22 develop the augmented quantile regression (AQR) estimator and apply it to a first-price auction model.
In this paper, we use an AQR estimator to estimate the distribution of bidder valuations. Our quantile estimator is the AQR counterpart to the estimator used in Gimenes, (2017). An AQR estimator is defined as the maximizer of an integrated version of the classic Koenker and Bassett, (1978) objective function, combined with kernel smoothing and polynomial approximation of the quantile function. Similar to the local polynomial estimator of Fan, (1992), advantages the AQR method has over standard QR are that it automatically yields estimators of both the quantile function and its derivative, and produces estimators with improved statistical properties near the boundaries. These features are crucial for our estimation problem, as the quantile derivative enters the FOC and all quantile levels of the AQR estimator can be used to estimate the risk-aversion parameter. By contrast, Gimenes, (2017) does not consider the FOC and therefore does not require estimation of the nuisance functions that enter it.
Other than the fact that the auction-specific relationship between quantiles of bidder valuations and bids differs between ascending auctions and first-price auctions, our application of the AQR method differs from GG22’s analysis in non-trivial ways due to our goal of estimating the risk-aversion parameter. In particular, we view the following two aspects as secondary contributions of the paper, as they expand the scope of the AQR method beyond our application. First, we construct and derive the asymptotic properties of a conditional CDF estimator defined as the inverse of an AQR estimator—see Lemma 3—which enters the FOC and does not appear in GG22. This estimator may be of independent interest as a flexible and low-dimensional approach to estimating conditional CDFs (see, e.g., Koenker et al., (2013) for a discussion on using a quantile function to estimate CDF and vice versa). Second, we show AQR estimators can be used as preliminary estimators in a two-step semiparametric M-estimation problem. Specifically, our risk-aversion parameter is an estimand defined implicitly as a solution to the FOC, whereas the finite-dimensional estimands considered in GG22 are explicit functionals of quantile objects. [^2] They motivate the study of such functionals with economic quantities that have closed-form expressions in the quantile functions (e.g., expected revenue (Li et al., (2003)) and bidder risk-aversion parameter in a first-price auction (Guerre et al., (2009))). Theorem 4 in GG22 shows that functionals of AQR estimators satisfy a central limit theorem (CLT). Our estimator does not take this form, but we show that it does so after suitable application of stochastic equicontinuity and linearization arguments, and that the CLT applies. This line of arguments is not new for developing limit theorems for semiparametric M-estimators. The novelty here is that we derive the linearized functional explicitly with AQR estimators as preliminary estimators, which has not been done previously. The econometrics and statistics literature is otherwise familiar with similar derivations when traditional nonparametric estimators, such as kernel and series estimators, are used—see, for example, Andrews, (1994), Newey, (1994), Chen et al., (2003), and Ichimura and Lee, (2010).
We perform a Monte Carlo study and find that our estimator behaves as expected from the theory. We then apply our methodology to real estate foreclosure auction data in São Paulo as an illustration. In this application, the sellers are lenders (banks and private companies) and the Court of Justice of the State of São Paulo. Possible reasons why these sellers may be risk averse include downside risks associated with failure to sell. For example, there may be future costs of holding such properties (e.g., property taxes, insurance, and security), the value of foreclosure homes may depreciate over time, and sellers may be subject to directives to sell within a certain time frame. [^3] Foreclosure assets are sometimes viewed as non-performing loans from an accounting perspective, and some central banks provide some guidelines on selling them accordingly. For example, see this ECB report: https://www.bankingsupervision.europa.eu/ecb/pub/pdf/guidance_on_npl.en.pdf.
Our empirical results suggest that sellers are risk-averse, and the model with a risk-averse seller fits the data much better than one with a risk-neutral seller. Notably, we find the model with risk-neutral sellers systematically overpredicts observed reserve prices relative to the model with risk-averse sellers. This finding complements several empirical evidence showing that observed reserve prices are well below the levels implied by the risk-neutral benchmark in a variety of auction settings (see McAfee and Vincent, (1992), Paarsch, (1997), McAfee et al., (2002), Haile and Tamer, (2003), and Tang, (2011)). Indeed, seller risk aversion has been proposed in the theoretical literature as one of the mechanisms that can rationalize these discrepancies (Maskin and Riley, (1984), Matthews, (1987), Hu et al., (2010)).
@@
We end the introduction with a review of related literature on the econometrics of auctions. The rest of the paper is organized as follows. Section 2 introduces the model and establishes key identification results for bidder and seller primitives. Section 3 develops the estimation methodology and the asymptotic properties of the AQR estimators for observed bids and bidder valuations. Section 4 presents the estimation of the risk-aversion parameter and derives its asymptotic properties. Section 5 reports a Monte Carlo study of the estimators. Section 6 provides an empirical application to real estate auctions in São Paulo. Section 7 concludes. Proofs of results are collected in the Appendix.
Background Literature
While there appear to be no prior econometric research on seller risk aversion, there is a growing literature studying risk-averse bidders, particularly in first-price auctions; see, for example Lu and Perrigne, (2008), Guerre et al., (2009), Campo et al., (2011), Li et al., (2015), Zincenko, (2018), Grundl and Zhu, (2019), and Jun and Zincenko, (2022). On the other hand, the empirical auction literature is built on the identification and inference of bidder valuation distributions under bidder risk neutrality. This begins with the parametric model of Paarsch, (1992) and grows with Donald and Paarsch (1993,1996), Laffont et al., (1995), Athey and Levin, (2001) amongst others.
Parametric models, however, can be computationally demanding to estimate and are subject to model misspecification. Nonparametric approaches, beginning with the seminal work of Guerre et al., (2000), provide an alternative and have since been developed in a variety of contexts; see, inter alia, Athey and Haile, (2002), Lu and Perrigne, (2008), Krasnokutskaya, (2011), Marmer and Shneyerov, (2012), Campo et al., (2011), Marmer et al., (2013), Enache and Florens, (2017), Liu and Luo, (2017), Luo and Wan, (2018) and Ma et al., (2019).
Purely nonparametric approaches are also not without limitations, as the convergence of nonparametric estimators deteriorates with the number of conditioning variables. Semiparametric approaches that aim to allow for more flexible modeling than parametric ones, without suffering from the curse of dimensionality, have been developed for mean regression (Rezende, (2008)) and quantile regression (Gimenes, (2017), Gimenes and Guerre, (2022)).
@@ 2 Model and identification
Consider an ascending auction of an indivisible object with $I\geq 2$ bidders. Bidders can raise prices continuously and without cost until only one bidder remains. The object is sold to the highest bidder for the price of his last bid, provided that it is at least as high as the reservation price. We assume an independent private values (IPV) environment, where each bidder knows only their own valuation and values are independently drawn across bidders.
Let the auctioned object have observable characteristics $X\in\mathbb{R}^{D}$. Bidder $i$’s valuation of the object is $V_{i}$, taking value in $\mathcal{V}=\left[\underline{v},\overline{v}\right]$. For notational simplicity, we assume $\mathcal{V}$ to be independent of $X$. We denote the conditional CDF of $V_{i}$ by $F\left(\cdot|X\right)$. To facilitate readers, as it may be instructive to compare our assumptions and results with GG22, we use the same notations and terminologies as them when possible.
@@ 2.1 Bidders’ behavior
The winning bid, denoted by $B=V^{I-1:I}\in\mathcal{V}$, is the $\left(I-1\right)$-th order statistic among the $I$ i.i.d. private values $\left\{V_{i}\right\}_{i=1}^{I}$. This is the same equilibrium play for ascending auctions as used in Aradillas-López et al., (2013) and Gimenes, (2017). We denote the conditional CDF of $B$ by $G\left(\cdot|X\right)$. We impose the relations between these variables in the following assumption.
Assumption M1.
(i) $F\left(\cdot|X\right)$ is continuous and strictly increasing almost surely on $\mathcal{V}$;
@@
We will take a quantile approach to model the bidder’s private value distribution. Let the $\alpha$-quantile of the valuation distribution be denoted by $V\left(\alpha|X\right)=F^{-1}\left(\alpha|X\right)$ for $\alpha\in\left[0,1\right]$. As done in Gimenes, (2017) and GG22, we take $V_{i}=V\left(A_{i}|X\right)$ where $A_{i}\sim Uni\left[0,1\right]$ can be viewed as the private rank of the bidder, which is independent of $X$ and other bidders’ ranks.
@@
Proposition 1 shows that the quantile functions of the winning bid and the private valuation are linked through a one-to-one relationship given in equation (1). This relation is the central identification argument on the bidder’s side (Gimenes, (2017)).
@@ 2.2 Seller’s behavior
We denote the seller’s value of the sale object by $W$. The seller can set the reserve price optimally to maximize her expected utility. Specifically, if the seller has utility function $U\left(\cdot\right)$, the expected utility from setting reserve price to be $r\in\mathcal{V}$ is:
@@
We denote the optimal reserve price by $R$. I.e., $R=\arg\max_{{}_{r\in\mathcal{V}}}\widetilde{\Pi}\left(r,X,W\right)$, whose existence and uniqueness are guaranteed under the conditions of Assumption M2 given below.
Analogously to writing the bidders’ bids in terms of private ranks, the expected revenue above can be equivalently expressed as a function of the screening level instead of the reserve price. For a screening level $\alpha$ that takes value in $\left[0,1\right]$, let us define
@@ 3 Augmented quantile regression
Given data on winning bids and auction characteristics, this section proposes estimators for the quantile function of private values and related functions under the linear quantile specification. Using Proposition 1, these estimators are obtained through appropriate transformations of the bid quantile function, which we estimate using the AQR method of GG22.
Abstracting from the auction interpretation, our AQR estimator for bids can also be viewed as a general-purpose quantile regression estimator, whose inverse provides an estimator of the conditional CDF. We establish pointwise asymptotic properties for the AQR estimators of these functions, as well as their uniform convergence rates, in Lemmas 1 and 3, respectively.
The statistical properties of the estimators for the distribution of private values are stated as propositions. We note that only convergence rates for these estimators are required to derive the large-sample properties of the risk-aversion estimator in Section 4. Moreover, we use only bid data in this section, as seller-specific variables—namely the reserve price and seller’s value of the auctioned object—are not used in the quantile estimation.